# Improved Margin Generalization Bounds for Voting Classifiers

### v1

@@ 5 Final Upper bound
$\left(\left\lfloor\frac{m}{\ln{\left(N_{i}\right)}}\right\rfloor+1\right)\delta_{i}\sup_{X\in{\mathcal{X}}^{2m}}{\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2})\leq\frac{m}{\ln{\left(N_{i}\right)}}\frac{\delta}{820N_{i}}\frac{d}{m(\gamma_{0}^{i})^{2}}.$ (45)
@@
Furthermore, we notice that by $N_{i}=\exp{\left(\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)},$ we have that
$\frac{m}{\ln{\left(N_{i}\right)}}\leq\frac{m(\gamma_{0}^{i})^{2}}{{72c^{\prime}d}\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}}\leq\frac{m(\gamma_{0}^{i})^{2}}{d},$ (46)
@@
where the last inequality follows by $c^{\prime}\geq 1$ and $\operatorname{\Gamma}{(\frac{16m(\gamma_{0}^{i})^{2}}{cd})}\geq 1$. Now, combining Section 5, Eq. 45 and Eq. 46 we get that
@@
$\leq\frac{\delta}{820N_{i}}=\frac{\delta}{820}\exp{\left(-\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}.$
We showed the above for any $i\in I=\{0,\ldots,\lfloor\log_{2}(1/\gamma^{\prime})\rfloor\}$; thus, by an application of the union bound, and $\gamma_{0}^{i}=\sqrt{2^{i}\gamma^{\prime}},$ we conclude that:
@@
$\leq\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\frac{\delta}{820}\exp{\left(-\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-\frac{72c^{\prime}d}{2^{i}\gamma^{\prime}}\right)}$
$\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-\frac{72c^{\prime}d}{2^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor-i}\gamma^{\prime}}\right)}\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-\frac{72c^{\prime}d2^{i}}{2^{\log_{2}(1/\gamma^{\prime})}\gamma^{\prime}}\right)}$
$\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-72c^{\prime}d2^{i}\right)}\leq\frac{\delta}{820},$
where the first inequality first inequality follows from the union bound, the second inequality by $\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}\geq 1,$ the third inequality by summing in the reverse order, the fourth inequality by $2^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\leq 2^{\log_{2}(1/\gamma^{\prime})},$ so $-1/2^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\leq-1/2^{\log_{2}(1/\gamma^{\prime})},$ and the last inequality by the sum being less than 1. Thus, we conclude that with probability at least $1-\delta$, we have that for all $i\in I$, for all $j\in J_{i}$, for all $\gamma\in\left[\gamma_{0}^{i},\gamma_{1}^{i}\right]$, for all $f\in\operatorname{\Delta({\mathcal{H}})}$, that either
@@
Furthermore, we have that $N_{i}=\exp{\left(\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}$, $\delta_{i}=\frac{\delta}{1640N_{i}^{2}}\frac{d}{m(\gamma_{0}^{i})^{2}}$, and by $\gamma\in\left[\gamma_{0}^{i},\gamma_{1}^{i}\right]$, with $\gamma_{1}^{i}\leq\sqrt{2}\gamma_{0}^{i}$, it implies that $\gamma\leq\sqrt{2}\gamma_{0}^{i}$, which combined gives that $\ln{\left(e/\delta_{i}\right)}$ can be bounded by
@@
$\leq\ln{\left(\frac{1640e}{\delta}\right)}+\frac{144c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{16m(\gamma_{0}^{i})^{2}}{cd}\right)}+\ln{\left(\frac{m\gamma^{2}}{d}\right)}$ (by $\gamma\geq\gamma_{0}^{i}$ )
$\leq\ln{\left(\frac{1640e}{\delta}\right)}+\frac{288c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}+\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}.$ (by $\gamma\leq\sqrt{2}\gamma_{0}^{i}$ and $\gamma_{0}^{i}\leq\gamma$ )
We notice that since $\ln{\left(x\right)}/x$ has derivative $\frac{1-\ln{\left(x\right)}}{x^{2}}=\frac{\ln{\left(e/x\right)}}{x^{2}}$ which is less than 00 for $x\geq e$, we have that $\operatorname{\mathrm{Ln}}{\left(x\right)}/x$ is decreasing for $x\geq e$ and since $\frac{1}{x}$ is decreasing for $0<x<e$ we conclude that $\operatorname{\mathrm{Ln}}\left(x\right)/x$ is decreasing for $x>0$. Thus, we conclude that
$\frac{\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}}{m}\leq\frac{d\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}}{m\gamma^{2}}\leq\frac{d\operatorname{\mathrm{Ln}}(\frac{m\gamma^{2}}{d})}{m\gamma^{2}}\leq\frac{d\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}}{m\gamma^{2}}.$
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Bigg{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{288c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}+\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}\right)}{m}}$
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{288c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}+\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}\right)}{m}\Bigg{)}$
$\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Bigg{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{289c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{289c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Bigg{)}.$
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Bigg{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{289c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$ (50)
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{289c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{16m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Bigg{)},$ (51)
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Big{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{9600\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$ (52)
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{9600\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Big{)}$
@@ 6 Majority of Three Large Margin Classifiers
To this end recall some notation. For a target concept $t\in\{-1,1\}^{{\mathcal{X}}}$ we use $({\mathcal{X}}\times\{-1,1\})_{t}^{*}$ for the set of all possible training sequences on ${\mathcal{X}}$, labelled by $t$, that is for $S\in({\mathcal{X}}\times\{-1,1\})_{t}^{*}$, any train example $(x,y)\in S$, is such that $y=t(x).$ We remark that $S$ is seen as a vector/sequence so it may have repetitions of similar train examples. Furthermore, for a distribution ${\mathcal{D}}$ over ${\mathcal{X}}$ we write ${\mathcal{D}}_{t}$ for the distribution over ${\mathcal{X}}\times\{-1,1\}$, defined by $\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}_{t}}\left[({\mathbf{x}},{\mathbf{y}})\in A\right]=\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[({\mathbf{x}},t({\mathbf{x}}))\in A\right]$ for any $A\subseteq{\mathcal{X}}\times\{-1,1\}$. Furthermore, for $R\subset{\mathcal{X}}$, such that $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R\right]\not=0$ , we define ${\mathcal{D}}_{t}\mid R$ as
$\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}_{t}\mid R}\left[({\mathbf{x}},{\mathbf{y}})\in A\right]=\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[({\mathbf{x}},t({\mathbf{x}}))\in A\mid{\mathbf{x}}\in R\right]=\frac{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[({\mathbf{x}},t({\mathbf{x}}))\in A,{\mathbf{x}}\in R\right]}{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R\right]}.$
We define a learning algorithm $L$ as a mapping from $({\mathcal{X}}\times\{-1,1\})_{t}^{*}$ to $\mathbb{R}^{{\mathcal{X}}}$, for $S\in({\mathcal{X}}\times\{-1,1\})_{t}^{*}$, we write $L_{S}$ for short of $L(S)\in\mathbb{R}^{{\mathcal{X}}}$. Furthermore, if $L_{S}\in\operatorname{\Delta({\mathcal{H}})}$ for any $S\in({\mathcal{X}}\times\{-1,1\})_{t}^{*}$ we write $L\in\operatorname{\Delta({\mathcal{H}})}$ For $10<\gamma<$ 1 and target concept $t$ we define a $\gamma$-margin algorithm $L$ for $t$ as a mapping from $({\mathcal{X}}\times\{-1,1\})_{t}^{*}$ to $\mathbb{R}^{{\mathcal{X}}}$, such that for a $S\in({\mathcal{X}}\times\{-1,1\})_{t}^{*}$, we have that $L_{S}(x)y\geq\gamma$ for all $(x,y)\in S.$ Furthermore, for three functions $f_{1},f_{2},f_{3}$ we define $\operatorname{\mathrm{Maj}}(f_{1},f_{2},f_{3})=\operatorname{\mathrm{sign}}(\operatorname{\mathrm{sign}}(f_{1})+\operatorname{\mathrm{sign}}(f_{2})+\operatorname{\mathrm{sign}}(f_{3}))$, with $\operatorname{\mathrm{sign}}(0)=0.$
@@
For distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $t$, hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ with VC-dimension $d$, training sequence size $m$, margin $10<\gamma<$ 1 and i.i.d. training sequences ${\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{t}^{m}$, it holds for any $\gamma$-margin learning algorithm $L\in\operatorname{\Delta({\mathcal{H}})}$ for $t$ that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(L_{{\mathbf{S}}_{1}},L_{{\mathbf{S}}_{2}},L_{{\mathbf{S}}_{3}}))\right]=O\left(\frac{d}{\gamma^{2}m}\right).$
Now as AdaBoost is a $\Omega(\gamma)$-margin learning algorithm when given access to a $\gamma$-weak learner ${\mathcal{W}},$ the above Theorem 11 implies Corollary 3.
@@
There exists a universal constant $C\geq 1$ such that: For distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $t$, hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ with VC-dimension $d$, training sequence size $m$, margin $10<\gamma<$ 1, i.i.d. training sequences ${\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{t}^{m}$, it holds for any $\gamma$-margin learning algorithm $L\in\operatorname{\Delta({\mathcal{H}})}$ for $t$ that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2}\sim{\mathcal{D}}^{m}_{t}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{1}}({\mathbf{x}}))\not=t({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{2}}({\mathbf{x}}))\not=t({\mathbf{x}})\right]\right]=\frac{96Cd}{\gamma^{2}m}.$

### v2

@@ 5 Final Upper bound
$\frac{2m}{\ln{\left(N_{i}\right)}}\delta_{i}\sup_{X\in{\mathcal{X}}^{2m}}{\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}^{i}\right\rceil},\frac{\gamma_{0}^{i}}{2})\leq\frac{m}{\ln{\left(N_{i}\right)}}\frac{\delta}{820N_{i}}\frac{d}{m(\gamma_{0}^{i})^{2}}.$ (45)
@@
Furthermore, we notice that by $N_{i}=\exp{\left(\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)},$ we have that
$\frac{m}{\ln{\left(N_{i}\right)}}\leq\frac{m(\gamma_{0}^{i})^{2}}{{72c^{\prime}d}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}}\leq\frac{m(\gamma_{0}^{i})^{2}}{d},$ (46)
@@
where the last inequality follows by $c^{\prime}\geq 1$ and $\operatorname{\Gamma}{(\frac{32m(\gamma_{0}^{i})^{2}}{cd})}\geq 1$. Now, combining Section 5, Eq. 45 and Eq. 46 we get that
@@
$\leq\frac{\delta}{820N_{i}}=\frac{\delta}{820}\exp{\left(-\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}.$
We showed the above for any $i\in I=\{0,\ldots,\lfloor\log_{2}(1/\gamma^{\prime})\rfloor\}$; thus, by an application of the union bound, and $\gamma_{0}^{i}=2^{i}\gamma^{\prime},$ we conclude that:
@@
$\leq\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\frac{\delta}{820}\exp{\left(-\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-\frac{72c^{\prime}d}{2^{2i}\gamma^{\prime 2}}\right)}$
$\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-\frac{72c^{\prime}d}{2^{2(\lfloor\log_{2}(1/\gamma^{\prime})\rfloor-i)}\gamma^{\prime 2}}\right)}\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-\frac{72c^{\prime}d2^{2i}}{2^{2\log_{2}(1/\gamma^{\prime})}\gamma^{\prime 2}}\right)}$
$\leq\frac{\delta}{820}\sum_{i=0}^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\exp{\left(-72c^{\prime}d2^{2i}\right)}\leq\frac{\delta}{820},$
where the first inequality first inequality follows from the union bound, the second inequality by $\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}\geq 1,$ the third inequality by summing in the reverse order, the fourth inequality by $2^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\leq 2^{\log_{2}(1/\gamma^{\prime})},$ so $-1/2^{\lfloor\log_{2}(1/\gamma^{\prime})\rfloor}\leq-1/2^{\log_{2}(1/\gamma^{\prime})},$ and the last inequality by the sum being less than 1. Thus, we conclude that with probability at least $1-\delta$, we have that for all $i\in I$, for all $j\in J_{i}$, for all $\gamma\in\left[\gamma_{0}^{i},\gamma_{1}^{i}\right]$, for all $f\in\operatorname{\Delta({\mathcal{H}})}$, that either
@@
Furthermore, we have that $N_{i}=\exp{\left(\frac{72c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}\right)}$, $\delta_{i}=\frac{\delta}{1640N_{i}^{2}}\frac{d}{m(\gamma_{0}^{i})^{2}}$, and by $\gamma\in\left[\gamma_{0}^{i},\gamma_{1}^{i}\right]$, with $\gamma_{1}^{i}\leq 2\gamma_{0}^{i}$, it implies that $\gamma\leq 2\gamma_{0}^{i}$, which combined gives that $\ln{\left(e/\delta_{i}\right)}$ can be bounded by
@@
$\leq\ln{\left(\frac{1640e}{\delta}\right)}+\frac{144c^{\prime}d}{(\gamma_{0}^{i})^{2}}\operatorname{\Gamma}{\left(\frac{32m(\gamma_{0}^{i})^{2}}{cd}\right)}+\ln{\left(\frac{m\gamma^{2}}{d}\right)}$ (by $\gamma\geq\gamma_{0}^{i}$ )
$\leq\ln{\left(\frac{1640e}{\delta}\right)}+\frac{576c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}+\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}.$ (by $\gamma\leq 2\gamma_{0}^{i}$ and $\gamma_{0}^{i}\leq\gamma$ )
Furthermore by $c^{\prime}\geq 1$ , we conclude that
$\frac{\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}}{m}\leq\frac{c^{\prime}d\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}}{m\gamma^{2}}\leq\frac{c^{\prime}d\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}}{m\gamma^{2}}.$
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Bigg{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{576c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}+\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}\right)}{m}}$
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{576c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}+\operatorname{\mathrm{Ln}}{\left(\frac{m\gamma^{2}}{d}\right)}\right)}{m}\Bigg{)}$
$\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Bigg{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{577c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{577c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Bigg{)}.$
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Bigg{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{577c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$ (50)
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{577c^{\prime}d}{\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Bigg{)},$ (51)
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)+64\Big{(}\sqrt{\frac{\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}(f)\cdot 2\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{20000\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}}$ (52)
$+\frac{5\left(\ln{\left(\frac{1640e}{\delta}\right)}+\frac{20000\tilde{c}^{2}\min(c,1)d}{32\gamma^{2}}\operatorname{\Gamma}{\left(\frac{32m\gamma^{2}}{\min(c,1)d}\right)}\right)}{m}\Big{)}$
@@ 6 Majority of Three Large Margin Classifiers
To this end we recall some notation. For a target concept $c^{*}\in\{-1,1\}^{{\mathcal{X}}}$ we use $({\mathcal{X}}\times\{-1,1\})_{c^{*}}^{*}$ for the set of all possible training sequences on ${\mathcal{X}}$, labelled by $c^{*}$, that is for $S\in({\mathcal{X}}\times\{-1,1\})_{c^{*}}^{*}$, any train example $(x,y)\in S$, is such that $y=c^{*}(x).$ We remark that $S$ is seen as a vector/sequence so it may have repetitions of similar train examples. Furthermore, for a distribution ${\mathcal{D}}$ over ${\mathcal{X}}$ we write ${\mathcal{D}}_{c^{*}}$ for the distribution over ${\mathcal{X}}\times\{-1,1\}$, defined by $\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}_{c^{*}}}\left[({\mathbf{x}},{\mathbf{y}})\in A\right]=\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[({\mathbf{x}},c^{*}({\mathbf{x}}))\in A\right]$ for any $A\subseteq{\mathcal{X}}\times\{-1,1\}$. Furthermore, for $R\subset{\mathcal{X}}$, such that $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R\right]\not=0$ , we define ${\mathcal{D}}_{c^{*}}\mid R$ as
$\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}_{c^{*}}\mid R}\left[({\mathbf{x}},{\mathbf{y}})\in A\right]=\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[({\mathbf{x}},c^{*}({\mathbf{x}}))\in A\mid{\mathbf{x}}\in R\right]=\frac{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[({\mathbf{x}},c^{*}({\mathbf{x}}))\in A,{\mathbf{x}}\in R\right]}{\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[{\mathbf{x}}\in R\right]}.$
We define a learning algorithm $L$ as a mapping from $({\mathcal{X}}\times\{-1,1\})_{c^{*}}^{*}$ to $\mathbb{R}^{{\mathcal{X}}}$, for $S\in({\mathcal{X}}\times\{-1,1\})_{c^{*}}^{*}$, we write $L_{S}$ for short of $L(S)\in\mathbb{R}^{{\mathcal{X}}}$. Furthermore, if $L_{S}\in\operatorname{\Delta({\mathcal{H}})}$ for any $S\in({\mathcal{X}}\times\{-1,1\})_{c^{*}}^{*}$ we write $L\in\operatorname{\Delta({\mathcal{H}})}$ For $10<\gamma<$ 1 and target concept $c^{*}$ we define a $\gamma$-margin algorithm $L$ for $c^{*}$ as a mapping from $({\mathcal{X}}\times\{-1,1\})_{c^{*}}^{*}$ to $\mathbb{R}^{{\mathcal{X}}}$, such that for a $S\in({\mathcal{X}}\times\{-1,1\})_{c^{*}}^{*}$, we have that $L_{S}(x)y\geq\gamma$ for all $(x,y)\in S.$ Furthermore, for three functions $f_{1},f_{2},f_{3}$ we define $\operatorname{\mathrm{Maj}}(f_{1},f_{2},f_{3})=\operatorname{\mathrm{sign}}(\operatorname{\mathrm{sign}}(f_{1})+\operatorname{\mathrm{sign}}(f_{2})+\operatorname{\mathrm{sign}}(f_{3}))$, with $\operatorname{\mathrm{sign}}(0)=0.$
@@
For distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $c^{*}$, hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ with VC-dimension $d$, training sequence size $m$, margin $10<\gamma<$ 1 and i.i.d. training sequences ${\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{c^{*}}^{m}$, it holds for any $\gamma$-margin learning algorithm $L\in\operatorname{\Delta({\mathcal{H}})}$ for $c^{*}$ that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(L_{{\mathbf{S}}_{1}},L_{{\mathbf{S}}_{2}},L_{{\mathbf{S}}_{3}}))\right]=O\left(\frac{d}{\gamma^{2}m}\right).$
Now as AdaBoost is a $\Omega(\gamma)$-margin learning algorithm when given access to a empirical $\gamma$-weak learner ${\mathcal{W}},$ the above Theorem 11 implies Corollary 3.
@@
There exists a universal constant $C\geq 1$ such that: For distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $c^{*}$, hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ with VC-dimension $d$, training sequence size $m$, margin $10<\gamma<$ 1, i.i.d. training sequences ${\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{c^{*}}^{m}$, it holds for any $\gamma$-margin learning algorithm $L\in\operatorname{\Delta({\mathcal{H}})}$ for $c^{*}$ that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2}\sim{\mathcal{D}}^{m}_{c^{*}}}\left[\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}\left[\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{1}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}}),\operatorname{\mathrm{sign}}(L_{{\mathbf{S}}_{2}}({\mathbf{x}}))\not=c^{*}({\mathbf{x}})\right]\right]=\frac{96Cd}{\gamma^{2}m}.$